OpenStudy (anonymous):

What is the general solution, to the equation (4cos^2 2x+1)sin(1/3)x=0? a)2pi n where n can be any real integer b)no solution c)3pi n where n can be any real integer d)(pi/3)n where n can be any real integer

OpenStudy (anonymous):

is it \[4\cos^{2}(2x+1)*\sin(x/3)= 0\] ?